OpenStudy (anonymous):

what is the solution of the system equation? -2x+y=3 8x-4y=-12

OpenStudy (anonymous):

First solve for y in the first equation

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

yeah its y=-2x+3

OpenStudy (anonymous):

Watch your negative sign

OpenStudy (anonymous):

Do you want to use elimination method or substitution method to solve this problem?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

Okay, retry to solve the first equation for y.

OpenStudy (anonymous):

idk cause thats how i thought u do it

OpenStudy (anonymous):

-2x + y = 3 +2x -2x + y = 3 + 2x y = 3 + 2x

OpenStudy (anonymous):

are u sure??

OpenStudy (anonymous):

You have to reverse the sign

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay nvm i see how u did it

OpenStudy (anonymous):

now we will substitute this into the second equation

OpenStudy (anonymous):

8x - 4(3 + 2x) = -12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you solve for x by yourself now?

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

There is something wrong. Double check your problem and make sure you typed the original correctly 8x - 12 - 8x = -12 0=0 The x term drops out completely, this shouldn't happen

OpenStudy (anonymous):

yeah i guess so thats why i ased cause i was getting confused

OpenStudy (anonymous):

I want to go back to this equation 8x - 12 - 8x = -12 rewritten as 8x - 12 = 8x - 12 lets get rid of the -12 8x = 8x then the 8's x=x So we can see that x equals all real numbers

OpenStudy (anonymous):

This question is written badly. You are positive this is the correct question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm going to admit that I am confused as to our answer as well. Because x can equal anything which gives us an infinite amount of solutions. x=y/2 - 3/2 8(y/2 - 3/2) - 4y = -12 4y -12 - 4y = -12 y can be anything as well I'd say there is any infinite amount of solutions

OpenStudy (anonymous):

This is a weird problem. Sry that I couln't help get an exact answer. All of our steps our correct, we didn't make any mistakes